OpenStudy (anonymous):

How do i rewrite this in ration exponent notation 3√7 (the 3 is supposed to be in the root)

myininaya (myininaya):

\[\sqrt[a]{x^b}=x^\frac{b}{a}\]

myininaya (myininaya):

\[\text{ so if } b=1 \text{ then } \sqrt[a]{x}=x^\frac{1}{a}\]

OpenStudy (anonymous):

ah thanks :D

myininaya (myininaya):

np